and overhead costs of $94,500 were estimated An 21,000 Number of inspections of Number of setups
Answer #1

Answer:

(a)

Determine the overhead rate for each activity.

Materials handling: $50,000 / 1000= $50.

Machine setup: $21,000 / 500 = $42.

Quality inspections: $33,000 / 750 = $44

__________________________________________________

(b)

Assign the manufacturing overhead costs for April to the two products using activity-based costing

Instruments

Activity

Rate

Total
Cost

A

B

C=A*B

Requisition

400

50

20000

Setups

200

42

8400

Inspection

200

44

8800

Total

37200

Per unit overhead cost: $37,200 / 50

744

Gauges

Activity

Rate

Total
Cost

A

B

C=A*B

Requisition

600

50

30000

Setups

300

42

12600

Inspection

400

44

17600

Total

60200

Per unit overhead cost: $60,200 / 300

201